Consider the stock of Aviva PLC. The beta of the stock is 1.25. The earnings per share (EPS) in the coming year is 55p. The expected Return on Equity (ROE) is 11%. The expected return on the market portfolio is 9%. The risk-free rate is 5%. The dividend pay-out ratio is 50%. What is the fair price of Aviva stock? 


1
Expert's answer
2021-02-04T14:44:10-0500

Solution:

The Cost of Capital (k) = "Risk\; free\; rate + Beta \;coefficient \times (Expected\; market\; return - Risk\; free \;rate)"


k = "0.05 + 1.25 \times (0.09 - 0.05)" 


Derive the dividends per share value:


Dividend payout ratio ="\frac{DPS}{EPS}"


"0.05 = \frac{DPS}{55}"


DPS ="0.05\times 55 = 2.75"


Dividend growth rate (g) = "(1 - Dividend\; Payout\; Ratio)\times ROE"

g ="(1 - 0.5)\times 0.11 = 0.5 \times0.11 = 0.055"


Fair value of stock ="\frac{DPS}{Required \;return\; of \;return\; (k) - Dividend\; growth\; rate\; (g)}"


FVS = "\frac{275}{(0.1-0.055) }"


FVS = "\frac{275}{0.45} = 61"


Fair Value of Stock = 61